OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 2. Find x when y = 4

OpenStudy (anonymous):

3xy=k 3(6)(2)=k 18(2)=k 36=k 3x(4)=36 12x=36 12x/12=36/12 x=3

OpenStudy (anonymous):

Isn't that direct variation?